12123333333333333333:

The function f(x) = 2x + 210 represents the number of calories burned when exercising, where x is the number of hours spent exercising. The function g(x) = 2x + 125 represents the calorie deficit that occurs when following a particular diet, where x is the number of hours spent exercising. What is (f + g)(2)? Explain. 214 calories burned while exercising for 2 hours. 214 calorie calories burned when combining diet with 2 hours of exercise. 343 calories burned while dieting for 2 hours. 343 calories burned while combining diet with 2 hours of exercise.

justus:

All we need to do is solve for (f + g)(2) because we already know that f(x) = 2x + 210 represents the number of calories burned when exercising and g(x) = 2x + 125 represents the calorie deficit that occurs when following a particular diet

12123333333333333333:

?

justus:

Wait actually

12123333333333333333:

But that isn't an answer... ?? I am sorry. I am dumbb

justus:

No you don't have to say that

justus:

Just a second I messed up

justus:

This equation thing is not working

justus:

I need to ask someone brb

12123333333333333333:

343 is correct...mchen said.

12123333333333333333:

He said D

justus:

Oof I am trying to put it into the equation thing and I just solved it but its not working

12123333333333333333:

okey new question

justus:

f+g=2x+210+2x+125\\\\f+g=4x+335\\\\(f+g)(2)=4\times 2+335\\\\(f+g)(2)=8+335=343

justus:

This is what I am trying to do

12123333333333333333:

mchen said it's D

justus:

Hes right

mhchen:

\[f+g = 2x+210\]